22. Suppose a market basket of goods and services costs $440 in the base year and the consumer price index
(CPI) is currently 125. This indicates the price of the market basket of goods is now:
a. $575.
b. $445.
c. $550.
d. $525.
23. Suppose the consumer price index (CPI) stands at 350 this year. If the inflation rate is 7.1 percent, then
next year's CPI will equal:
a. 350.
b. 360.
c. 375.
d. 500.
Answer #1

22) The price of the market basket of goods is now $550

Consumer Price Index = Price of market basket / Price of market basket in base year x 100

125 = Price of market basket / 440 x 100

125 x 440 = Price of market basket x100

55000 = Price of market basket x 100

55000 / 100 = Price of market basket

Price of market basket = $550

Hence option C is correct

23) Next year's CPI will equal be 375

Inflation Rate = Next CPI - Current CPI / Current CPI x 100

7.1 = Next CPI - 350/ 350 x 100

7.1 x 350 = (Next CPI - 350) x 100

2485 + 35000 = Next CPI x 100

37485 / 100 = Next CPI

Next CPI = 374.85 or 375

Hence option C is correct